OpenStudy (anonymous):

A line passes through point (-2,5) and has a slope of 2/3. Points A(x,3) and B(-2,y) lie on the line. What's the value of x, and y?

myininaya (myininaya):

do you know how to find the slope of a line?

myininaya (myininaya):

I mean this: Say I give you two points on a line, how would you find the slope of that line?

myininaya (myininaya):

For example, if I say (-2,5) and (x,3) is on a line, then the slope of this line is?

myininaya (myininaya):

Here is the formula I'm asking you to use: \[\text{ Then slope of a line going through } (x_1,y_1) \text{ and } (x_2,y_2) \\ \text{ is } m=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

i think i figured it out.. value of x is 1 and value of y is 5?

myininaya (myininaya):

let me check

myininaya (myininaya):

\[\text{ The slope of a line going through } (-2,5) \text{ and } (x,3) \\ \text{ is } m=\frac{3-5}{x-(-2)}=\frac{-2}{x+2} \\ \text{ we are given } m=\frac{2}{3} \text{ though } \\ \frac{2}{3}=\frac{-2}{x+2} \\ \frac{2}{3}=\frac{2}{-(x+2)} \\ \implies 3=-(x+2)\]

myininaya (myininaya):

if you solve that equation you do not get x=1

myininaya (myininaya):

also your y value is right

myininaya (myininaya):

(-2,5) was given that it was already on the line so (-2,y) would mean y has to be 5 :)

myininaya (myininaya):

the only value that needs work is your x value

OpenStudy (anonymous):

okay thank you

myininaya (myininaya):

did you find the correct x value ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

someone please help! lol